OpenStudy (anonymous):

If w = <-8, 15>, find ||w||. 7 17 23 sqr7 sqr23

OpenStudy (zzr0ck3r):

\(w=\langle a, b\rangle \implies \ ||w|| = ||\langle a, b \rangle|| =\sqrt{a^2+b^2} \)

OpenStudy (anonymous):

so -8^2+15^2

OpenStudy (zzr0ck3r):

\(\sqrt{(-8)^2+15^2}=\)

OpenStudy (anonymous):

i got 12.6

OpenStudy (phi):

remember that -8 * -8 is +64 (not -64)